OpenStudy (anonymous):

What is the equation of the line passing through (0, 2) and perpendicular to the line y = 4x – 3 in slope-intercept form?

OpenStudy (anonymous):

im learning this right now too so i will tell you as much as i know

OpenStudy (anonymous):

but the slope would be opposite of 4 because its perpendicular so the slope for your new equation is going to be -1/4

OpenStudy (anonymous):

your new equation is y-2=-1/4(x-0) which when solved is y=-1/4x+2

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

yeah, thanks.

OpenStudy (anonymous):

m=-1/4 is the gradient of the new line so just use y-y1=m(x-x1) so y-2=-1/4(x-0)

OpenStudy (anonymous):

yeah and -1/4(0)=0

OpenStudy (anonymous):

so then you add 2 to both sides and get y=-1/4+2

OpenStudy (anonymous):

ohh yeh oops

OpenStudy (anonymous):

thats it